# The Power of Training: How Different Neural Network Setups Influence the Energy Demand

### v2

@@ 2 Related Works
Due to lacking practicability and the complexity of tasks, precise energy consumption tracking is still not commonly applied in the field of AI and Machine Learning [9]. However, the rising demand for energy must be counteracted to minimize the environmental footprint and economic expenses, for instance when hosting large server and HPC farms, in order to operate sustainably in the future [17]. In the past focus lay on inference optimization for edge devices, microcontrollers, and field-programmable gate arrays (FPGAs), since the energy demands of training machine learning models have received less attention due to smaller implications [6]. Nevertheless, with the rise of Large Language Models (LLM) and other complex architectures, the computational demands for training such models have a significantly larger share of the lifecycle energy requirement [20, 16]. This in turn requires a precise consideration of possible adjustment screws to counteract these trends, whereas this work focuses on the exploration and tuning of hyperparameters [10]. To the best of our knowledge, this work is the first to investigate the selection of popular hyperparameter spaces with a hardware-based focus on quantizing the energy demand instead of purely aiming towards optimal model performance.
@@ 2.2 Hyperparameter Selection
All strategies together share a downside: They purely focus on maximizing the model performance without incorporating the energy consumption [8]. This oversight leads to an enhanced waste of energy and usually leads to training and discarding models that don’t meet the required performance [25]. Therefore, there is a need to investigate the correlation between hyperparameter selection and energy demand in order to develop strategies that may strive for the best performance while maintaining moderate energy demands.
@@ 4 Training Regime Results
Figure 3: Energy per epoch across the different application scenarios with varying Batch Size and Learning Rate.
@@
The average energy that was demanded during the training is shown in Figure 3, calculated through the introduced equation Equation 1. As a general rule, the level of energy used can be linked to the efficiency and utilization of the GPU. Within each matrix, lower values mean less energy consumed for training one epoch, which results in improved GPU efficiency.
For each matrix shown in table Figure 3, the batch size choice has the most impact on the energy consumption. Since the batch size determines the level of operations that can be executed in parallel, effects in the energy per epoch mainly originate from the utilization and the efficiency window in which the GPU is operated. There needs to be a balance between the dataset size and the GPU capabilities in order to operate it efficiently. Cross-comparing the matrices across the two application scenarios highlights the importance of proper hyperparameter utilization. Due to the large size of the CelebA image dataset with each data point representing a 78×218 pixel-sized image, the smallest batch size can be considered the best selection to operate the GPU efficiently On the other hand, the tremendously smaller size of the PAMAP2 results in in-efficient utilization of the GPU for small batch sizes. Therefore, larger batch sizes run more efficiently for the PAMAP2 scenario since less energy is wasted for GPU idling.
@@
Overall, the range between the worst and best-performing setup is wide, with the best case usually consuming less than 20% of the worst-case setup. As a guideline to achieve this much of an improvement, the energy per epoch and necessary number of epochs both need to be minimized. Comparing our two application scenarios, especially when checking the CelebA results in Figure 4, there can not be derived a general rule for thoroughly choosing the best hyperparameters. Considering the extension with less prominent hyperparameters, for instance These findings therefore provide a motivation for further research to explore more dynamic and thorough hyperparameter selection approaches.
@@ 6 Multitask Learning Paradigm Results
The experiment was conducted through the same procedure as the previous pretraining scenario, training the three models (attribute, identity, and both together) with the 16 different setups. We calculated the average energy per epoch across the three scenarios as shown in Table 3. Since the average energy per epoch is almost identical across training attributes, identity, or both, with different operating levels across the three hardware configurations, we further calculated the average length of the training in number of epochs. As we can see, for this experiment the number of epochs rules the overall energy consumption whereas the efficiency from energy per epoch is negligible. Since the loss functions are connected during the training in the multitask scenario, the classifiers of attribute and identity are able to cross-regulate each other, which results in a quicker convergence. Instead of recycling already present knowledge as for the pretraining, this experiment benefits from the active exchange of information through the loss function. The result manifests our previous findings for minimizing either one or at best both, the energy per epoch and the overall number of epochs, to properly decrease the energy consumption for the machine learning training.
@@ 7 Correlation with High-End Consumer Hardware
For the training regime, we could reproduce the influences of batch size on the energy per epoch efficiency metric and the learning rate influence towards However, the findings were somewhat less distinctive and showed higher variations. We attribute this to the missing isolation of the jobs, which are now running on commercial operating systems. Additionally, for the powermetrics-based tracking, it remains unclear how precise the measurements can be compared to the carbontracker measurements. Even though the local hardware operated on different power levels, running the two learning paradigm tests confirmed our findings that pretraining and multitask learning can be utilized to improve life-cycle-oriented efficiency. For pretraining, the average breakeven points are calculated for 9 cycles on the CUDA setup and 12 cycles on MPS. Finally, for the multitask learning, the information sharing through the loss function by training attribute and identity at the same time could reduce the energy consumption by around 40%.

### v3

@@ 2 Related Works
Due to lacking practicability and the complexity of tasks, precise energy consumption tracking is still not commonly applied in the field of AI and Machine Learning [9]. However, the rising demand for energy must be counteracted to minimize the environmental footprint and economic expenses, for instance when hosting large server and HPC farms, in order to operate sustainably in the future [16]. In the past focus lay on inference optimization for edge devices, microcontrollers, and field-programmable gate arrays (FPGAs), since the energy demands of training machine learning models have received less attention due to smaller implications [6]. Nevertheless, with the rise of Large Language Models (LLM) and other complex architectures, the computational demands for training such models have a significantly larger share of the lifecycle energy requirement [19]. This in turn requires a precise consideration of possible adjustment screws to counteract these trends, whereas this work focuses on the exploration and tuning of hyperparameters [10]. To the best of our knowledge, this work is the first to investigate the selection of popular hyperparameter spaces with a hardware-based focus on quantifying the energy demand instead of purely aiming towards optimal model performance.
@@ 2.2 Hyperparameter Selection
All strategies together share a downside: They purely focus on maximizing the model performance without incorporating the energy consumption [8]. This oversight leads to an enhanced waste of energy and usually leads to training and discarding models that do not meet the required performance [24]. Therefore, there is a need to investigate the correlation between hyperparameter selection and energy demand in order to develop strategies that may strive for the best performance while maintaining moderate energy demands.
@@ 4 Training Regime Results
Figure 3: Energy per epoch (mWh) across the different application scenarios with varying Batch Size and Learning Rate.
@@
The average energy that was demanded during the training is shown in Figure 3, calculated through the introduced equation Equation 1. For this equation, k states the number of epochs in range 0 to T and n represents the number of power measurements within an epoch from which we calculate the average. As a general rule, the level of energy used can be linked to the efficiency and utilization of the GPU. Within each matrix, lower values mean less energy consumed for training one epoch, which results in improved GPU efficiency.
For each matrix shown in table Figure 3, the batch size choice has the most impact on the energy consumption. Since the batch size determines the level of operations that can be executed in parallel, effects in the energy per epoch mainly originate from the utilization and the efficiency window in which the GPU is operated. There needs to be a balance between the dataset size and the GPU capabilities in order to operate it efficiently. Cross-comparing the matrices across the two application scenarios highlights the importance of proper hyperparameter utilization. Due to the large size of the CelebA image dataset with each data point representing a 78×218 pixel-sized image, the smallest batch size can be considered the best selection to operate the GPU efficiently. On the other hand, the tremendously smaller size of the PAMAP2 results in in-efficient utilization of the GPU for small batch sizes. Therefore, larger batch sizes run more efficiently for the PAMAP2 scenario since less energy is wasted for GPU idling.
@@
Overall, the range between the worst and best-performing setup is wide, with the best case usually consuming less than 20% of the worst-case setup. As a guideline to achieve this much of an improvement, the energy per epoch and necessary number of epochs both need to be minimized. Comparing our two application scenarios, especially when checking the CelebA results in Figure 4, there can not be derived a general rule for thoroughly choosing the best hyperparameters. These findings therefore provide a motivation for further research to explore more dynamic and thorough hyperparameter selection approaches.
@@ 6 Multitask Learning Paradigm Results
The experiment was conducted through the same procedure as the previous pretraining scenario, training the three models (attribute, identity, and both together) with the 16 different setups. We calculated the average energy per epoch across the three scenarios as shown in Table 3. Since the average energy per epoch is almost identical across training attributes, identity, or both, with different operating levels across the three hardware configurations, we further calculated the average length of the training in number of epochs. As we can see, for this experiment the number of epochs rules the overall energy consumption whereas the efficiency from energy per epoch is negligible. Since the loss functions are connected during the training in the multitask scenario, the classifiers of attribute and identity are able to cross-regulate each other, which results in a quicker convergence. Instead of recycling already present knowledge as for the pretraining, this experiment benefits from the active exchange of information through the loss function. The result manifests our previous findings for minimizing either one or at best both, the energy per epoch and the overall number of epochs, to properly decrease the energy consumption for the training.
@@ 7 Correlation with High-End Consumer Hardware
For the training regime, we could reproduce the influences of batch size and the learning rate influence on the energy per epoch efficiency metric. However, the findings were somewhat less distinctive and showed higher variations. We attribute this to the missing isolation of the jobs, which are now running on commercial operating systems. Additionally, for the powermetrics-based tracking, it remains unclear how precise the measurements can be compared to the carbontracker measurements. Even though the local hardware operated on different power levels, running the two learning paradigm tests confirmed our findings that pretraining and multitask learning can be utilized to improve life-cycle-oriented efficiency. For pretraining, the average breakeven points are calculated for 9 cycles on the CUDA setup and 12 cycles on MPS. Finally, for the multitask learning, the information sharing through the loss function by training attribute and identity at the same time could reduce the energy consumption by around 40%.